OpenStudy (anonymous):

Determine all horizontal, slant, and vertical asymptotes. For each vertical asymptote, determine whether f(x) -> infinity sign of f(x) -> negative infinity sign on either side of the asymptote.

OpenStudy (anonymous):

f(x) = x^2/4 - x^2

OpenStudy (anonymous):

is it \[\frac{x^2}{4-x^2}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so for the vertical asymptotes, set the denominator equal to zero and solve for \(x\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you get \[4-x^2=0\] or \[(2-x)(2+x)=0\] and the solutions are \(x=-2\) or \(x=2\) those are your vertical asymptotes (there are two of them)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for the horizontal asymptote, note that the numerator and denominator have the same degree (both are degree 2) so it is the ratio of the leading coefficients

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the leading coefficient of \(x^2\) is 1 and the leading coefficient of \(4-x^2\) is \(-1\) therefore the horizontal asymptote is \(y=-1\)

OpenStudy (anonymous):

that is in the case where the degrees are the same. there is no slant asymptote for there to be a slant asymptote , the degree of the numerator would have to be one more than the degree of the denominator

OpenStudy (anonymous):

okay thx

OpenStudy (anonymous):

yw